OpenStudy (anonymous):

Solve by factoring. If a number is added to its square, the result is 56. Find the number.

OpenStudy (anonymous):

n + n^2 = 56 n^2+n-56=0 (n-7) (n+8)=0 n=7 or -8

OpenStudy (anonymous):

robtobey is correct